Two ships A and B sail simultaneously from the same place. vessel A on course 300at 15 knots and B on course N 45°E at 10 knots. After 1 hour and 30 min, how far is vessel A to vessel B?




1
Expert's answer
2022-05-10T08:13:30-0400

As theorem of cosine

"c^2=a^2+b^2-2abcos \alpha"

a=15x1.5=22.5 knocks

b=10x1.5=15 knocks

"c=\sqrt{22.5^2+15^2-2x22.5x15xcos (45°)}=\sqrt{506.25+225-477.3}=\sqrt{253.95}=15.9 knokcs"